# A Recipe of Parallel Corpora Exploitation for Multilingual Large Language Models

### v1

@@ Abstract
Recent studies have highlighted the potential of exploiting parallel corpora to enhance multilingual large language models, improving performance in both bilingual tasks, e.g., machine translation, and general-purpose tasks, e.g., text classification. Building upon these findings, our comprehensive study aims to identify the most effective strategies for leveraging parallel corpora. We investigate the impact of parallel corpora quality and quantity, training objectives, and model size on the performance of multilingual large language models enhanced with parallel corpora across diverse languages and tasks. Our analysis reveals several key insights: (i) filtering noisy translations is essential for effectively exploiting parallel corpora, while language identification and short sentence filtering have little effect; (ii) even a corpus containing just 10K parallel sentences can yield results comparable to those obtained from much larger datasets; (iii) employing only the machine translation objective yields the best results among various training objectives and their combinations; (iv) larger multilingual language models benefit more from parallel corpora than smaller models due to their stronger capacity for cross-task transfer. Our study offers valuable insights into the optimal utilization of parallel corpora to enhance multilingual large language models, extending the generalizability of previous findings from limited languages and tasks to a broader range of scenarios.
@@ 1 Introduction
Figure 1: Average performance improvements achieved by mLLMs enhanced with parallel corpora compared to their base models. Best: Instruction tuning of BLOOM-7B1 with the machine translation objective (MT) using 10K high-quality parallel sentences yields the best results. Main variations explored include: Filter (No) (using the original data); OBJ (TLM) (translation language modeling objective); OBJ (XSS) (cross-lingual semantic similarity objective); |Data| (50K) (a larger 50K-sentence dataset); |Model| (1B7) (BLOOM-1B7 model).
@@
However, existing studies often overlook a comprehensive exploration of methodologies for harnessing parallel corpora. The quality and quantity of parallel corpora remain inadequately explored, inhibiting the full potential of such resources. Moreover, the influence of different training objectives and mLLM sizes across diverse languages and tasks remains under-investigated. This limitation impedes the generalization of parallel corpora exploitation methods across varied linguistic landscapes and task domains. Therefore, this paper aims to address these gaps by presenting a comprehensive recipe for exploiting parallel corpora for mLLMs. We focus on four key factors, with some main results shown in Figure 1.
Quality: Kreutzer et al. (2022) highlight large portions of low-quality data within current massive parallel corpora. We explore three dimensions of quality: translation accuracy, sentence length, and language identification. Our findings emphasize the critical role of translation quality in exploiting parallel corpora, while showing that sentence length filtering and language identification have minimal impact.
Quantity: Acquiring substantial amounts of high-quality parallel corpora presents a significant challenge, especially for low-resource languages. Our study examines the minimum corpus size necessary to achieve performance improvements across diverse tasks. Remarkably, we find that even a corpus of just 10K sentences can yield results comparable to those obtained from much larger datasets.
Objective: Previous studies (Cahyawijaya et al., 2023) have investigated the effectiveness of different training objectives and their combinations on classification tasks of Indonesian local languages, using smaller-sized mLLMs up to 1B7. We extend this investigation by examining the impact of various training objectives and their combinations on larger mLLMs across a range of languages and tasks. Our experiments demonstrate that employing the machine translation objective produces the most promising results.
@@
In light of the critical role parallel corpora play in mLLMs, our study provides a comprehensive recipe for effectively exploiting parallel corpora. We have identified four primary factors: quality (§4), quantity (§5), objective (§6), and model size (§7). Our detailed analysis of these factors reveals their great impact on mLLM performance across diverse languages and tasks. By delving into these aspects, we offer actionable insights that can inform the development and optimization of strategies for parallel corpora exploitation, ultimately contributing to the advancement of mLLMs in both bilingual and general-purpose tasks.
@@ 2.1 Parallel Data for Multilingual Language Models
In the era of mLLMs, parallel corpora are constructed as instruction data and used to enhance mLLMs through supervised fine-tuning (Cahyawijaya et al., 2023; Zhu et al., 2023; Li et al., 2023). Specifically, Cahyawijaya et al. (2023) propose three methods of incorporating parallel corpora as instruction tuning data: Machine Translation (MT), Translation Language Modeling (TLM), Cross-Lingual Semantic Similarity (XSS) (see §3.3). However, their evaluation is limited to small models with up to 1.7 billion parameters and focuses solely on classification tasks within Indonesian local languages. Both Zhu et al. (2023) and Li et al. (2023) propose using machine-translation-style instruction data to improve mLLMs but do not explore different training objectives. While these studies yield promising results, their scope is limited. Firstly, they fail to explore critical factors such as the quality and quantity of parallel corpora, considering the high cost of collecting high-quality and massive parallel corpora, especially for low-resource languages. Secondly, their investigations do not encompass an in-depth analysis of training objectives and mLLMs with varied model sizes across diverse languages and tasks.
@@ 2.2 Key Elements for Language Modeling
Quality: Kreutzer et al. (2022) conducted manual audits of prevalent monolingual and parallel corpora, revealing significant portions of low-quality data, particularly in corpora for low-resource languages. Follow-up studies have investigated the impact of data quality on model performance. Artetxe et al. (2022) observed that similar results on downstream tasks can be achieved regardless of the degree of quality of the corpus used for pretraining, while other studies found that the quality of parallel corpora matters for machine translation (Ranathunga et al., 2024) and general-purpose tasks (Reid and Artetxe, 2022a).
@@
Model Size: Recent studies indicate that scaling up language models enhances their capability to excel in diverse and complex reasoning tasks (Wei et al., 2022, 2023; Lu et al., 2023). Follow-up studies (DBLP:conf/nips/ShuW0GXG23; Wei et al., 2023) further illustrate distinct behavioral differences between larger and smaller models.
@@ 3.1 Language
We use two criteria for language selection. Firstly, we select languages well covered by mLLMs and evaluation benchmarks, allowing for robust evaluation across diverse downstream tasks. Secondly, we select typologically diverse languages, enabling our investigation to generalize to a wide range of low-resource languages. Thus, we select five languages: Arabic (ar), Spanish (es), Hindi (hi), Vietnamese (vi) and Chinese (zh).
@@ 3.3 Training
We select the BLOOM series (Scao et al., 2022) for our investigation due to its offering of different sizes of mLLMs pretrained for the five target languages under consideration. We explore BLOOM models of various sizes, including 7B1, 3B, and 1B7. Due to limited computational resources, we use LoRA (Hu et al., 2022), which is known for its competitive performance compared to full-parameter training (Alves et al., 2023), for instruction tuning of BLOOM. We configure the learning rate to $41e-$ 4, weight decay to $0.1$, and set the rank of LoRA to $16$. The maximum sequence length for both source and target sentences is set to $128$. To maintain consistency across experiments with different quantities of parallel corpora, we ensure a uniform training budget of 50K parallel sentences. Specifically, we calculate the number of epochs as 50K divided by the number of sentences considered from the OPUS100 dataset. The batch size is $128$, and we save the checkpoints every 20 steps.
Following Cahyawijaya et al. (2023), we construct the data for instruction tuning based on the parallel corpora by three distinct patterns: Machine Translation (MT), Translation Language Modeling (TLM), and Cross-Lingual Semantic Similarity (XSS). Table 6 presents the templates for these three objectives. Here, [SOURCE_LANG] and [TARGET_LANG] represent the language names of the source and target languages, respectively. In our study, we consider both English-to-target-language and target-language-to-English directions, where [SOURCE_LANG] represents English or [TARGET_LANG] represents English. For MT, [SOURCE_TEXT] and [TARGET_TEXT] refer to the parallel sentences in the source and target languages, respectively. For TLM, a portion of tokens in [TARGET_TEXT] are masked to generate [INPUT_TEXT]. For XSS, our objective is to predict whether parallel sentences [SOURCE_TEXT] and [TARGET_TEXT] are semantically similar, with [LABEL] being “Yes” or “No”. Specifically, we utilize the parallel corpora as positive examples and introduce perturbations to [TARGET_TEXT] to construct negative examples. We consider applying the objectives both individually and in combination.
@@ 3.4 Evaluation
To provide a thorough understanding of our evaluation procedures, we offer detailed prompts for each task in §A. In all experiments, we employ a 2-shot in-context learning approach, where the model is given two examples appended to the input query to aid in making predictions.
@@ 4.1 Quality of OPUS100
Low-resource languages suffer more from low-quality issues. For low-resource languages like Hindi, there are fewer high-quality parallel sentences compared to high-resource languages such as Spanish. Analysis of translation quality indicates that the English-Hindi pair has less than 20% of parallel sentences with high COMETWIKI scores (0.75-1), whereas the English-Spanish pair has around 45%. For sentence length, the English-Hindi pair contains 10% more short sentences (0-5 tokens) compared to high-resource language pairs. Moreover, both the English-Arabic and English-Hindi pairs exhibit about 10% more parallel sentences that may be in the wrong languages.
@@ 4.2 Effect of Quality
Parallel corpora containing noisy translations still improve results. Comparing the results of the experiment with $\tau_{c}=0$ (ID 1) to the original model (ID 0), there’s an average improvement of 0.4% for all tasks. The most notable improvements are observed in both bilingual tasks (en-xx) and in-language tasks. However, generating English for bilingual tasks yields degraded or marginally improved results. Experiment 0 exhibits 0.7% and 1.1% decrements in FLORES and MUSE respectively, with only a 0.3% improvement in MLQA.
@@ 5.1 Effect of Quantity Across Tasks
Table 4: Performance (%) of BLOOM-7B1 after instruction tuning with the machine translation objective using varying amounts of parallel sentences, obtained with the best filtering strategy (ID 5) as shown in Table 3. |Sent| indicates the number of parallel sentences used for instruction tuning, with |Sent|=0 representing the original BLOOM-7B1 model.
@@
Using 10K parallel sentences leads to the optimal performance. The best overall performance is achieved with 10K parallel sentences, resulting in an average score of 54.0%. This setting yields the highest scores in MUSE and SIB.
@@
Table 5: Performance (%) of BLOOM-7B1 after instruction tuning with the machine translation objective using varying amounts of parallel sentences, obtained with the best filtering strategy (ID 5) as shown in Table 3. |Sent| indicates the number of parallel sentences used for instruction tuning, with |Sent|=0 representing the original BLOOM-7B1 model.
@@ 5.2 Effect of Quantity Across Languages
We delve deeper into the influence of parallel corpora quantity across various languages, as depicted in Table 5.
@@ 6 Objective
We explore the effectiveness of different objectives and their combinations, with results presented in Table 6.
@@ 8 Conclusion
Appendix A Prompt

### v2

@@ Abstract
Recent studies have highlighted the potential of exploiting parallel corpora to enhance multilingual large language models in both bilingual tasks, e.g., machine translation, and general-purpose tasks, e.g., text classification. Building upon these findings, our comprehensive study aims to identify the most effective strategies for leveraging parallel corpora. We investigate the impact of parallel corpus quality and quantity, training objectives, and model size on the performance of multilingual large language models enhanced with parallel corpora across diverse languages and tasks. Our analysis reveals several key insights: (i) filtering noisy translations is essential for exploiting parallel corpora, while language identification and short sentence filtering have little effect; (ii) even a corpus with just 10K parallel sentences can yield results comparable to those obtained from larger datasets; (iii) employing only the machine translation objective yields the best results among various training objectives and their combinations; (iv) larger multilingual language models benefit more from parallel corpora than smaller models. Our study offers valuable insights into the optimal utilization of parallel corpora to enhance multilingual large language models, extending the generalizability of previous findings from limited languages and tasks to a broader range of scenarios.
@@ 1 Introduction
Figure 1: Average performance improvements (y-axis) achieved by mLLMs enhanced with parallel corpora compared to their base models. Best: Instruction tuning of BLOOM-7B1 with the machine translation objective (MT) using 10K high-quality (i.e., filtered) parallel sentences yields the best results. Main variations explored include: Filter (No) (using the original data); OBJ (TLM) (translation language modeling objective); OBJ (XSS) (cross-lingual semantic similarity objective); |Data| (50K) (a larger 50K-sentence dataset); |Model| (1B7) (BLOOM-1B7 model).
@@
However, existing studies often lack in comprehensive exploration of methodologies for harnessing parallel corpora. The quality and quantity of parallel corpora remain inadequately explored, inhibiting the full potential of such resources. Moreover, the influence of different training objectives and mLLM sizes across diverse languages and tasks remains under-investigated. This limitation impedes the generalization of parallel corpus exploitation methods across varied linguistic landscapes and task domains. Therefore, this paper aims to address these gaps by presenting a comprehensive recipe for exploiting parallel corpora for mLLMs. We focus on four key factors, with some main results shown in Figure 1.
Quality: We explore three dimensions of parallel corpus quality: translation accuracy, sentence length, and language identification. Our results show that translation quality is vital for exploiting parallel corpora, while sentence length filtering and language identification have minimal impact.
Quantity: Acquiring large amounts of high-quality parallel corpora is challenging, especially for relatively low-resource languages. Our study examines the minimum corpus size necessary to achieve performance improvements across diverse tasks. We find that even a corpus of just 10K sentences can yield results comparable to those obtained from much larger datasets.
Objective: Previous studies (Cahyawijaya et al., 2023) have investigated the effectiveness of different training objectives and their combinations on classification tasks of Indonesian local languages, using smaller-sized mLLMs up to 1B7 parameters. We extend this investigation by examining the impact of various training objectives and their combinations on larger mLLMs across a range of languages and tasks. Our experiments demonstrate that employing the machine translation objective produces the most promising results.
@@
In light of the critical role parallel corpora play in mLLMs, our study provides a comprehensive recipe for effectively exploiting parallel corpora. We have identified four primary factors: quality (§4), quantity (§5), objective (§6), and model size (§7). Our detailed analysis of these factors reveals their great impact on mLLM performance across diverse languages and tasks. By delving into these aspects, we offer actionable insights that can inform the development and optimization of strategies for parallel corpus exploitation, ultimately contributing to the advancement of mLLMs in both bilingual and general-purpose tasks.
@@ 2.1 Parallel Data for Multilingual Language Models
In the era of mLLMs, parallel corpora are constructed as instruction data and used to enhance mLLMs through supervised fine-tuning (Cahyawijaya et al., 2023; Zhu et al., 2023; Li et al., 2023). Specifically, Cahyawijaya et al. (2023) propose three methods of incorporating parallel corpora as instruction tuning data: Machine Translation (MT), Translation Language Modeling (TLM), and Cross-Lingual Semantic Similarity (XSS) (see §3.3). However, their evaluation is limited to small models with up to 1.7 billion parameters and focuses solely on classification tasks within Indonesian local languages. Both Zhu et al. (2023) and Li et al. (2023) propose using machine-translation-style instruction data to improve mLLMs but do not explore different training objectives. While these studies yield promising results, their scope is limited. Firstly, they do not explore critical factors such as the quality and quantity of parallel corpora, considering the high cost of collecting high-quality and massive parallel corpora, especially for relatively low-resource languages. Secondly, their investigations do not encompass an in-depth analysis of training objectives and mLLMs with varied model sizes across diverse languages and tasks.
@@ 2.2 Key Elements for Language Modeling
Quality: Kreutzer et al. (2022) conducted manual audits of prevalent monolingual and parallel corpora, revealing significant portions of low-quality data, particularly in corpora for relatively low-resource languages. Follow-up studies have investigated the impact of data quality on model performance. Artetxe et al. (2022) observed that similar results on downstream tasks can be achieved regardless of the degree of quality of the corpus used for pretraining, while other studies found that the quality of parallel corpora matters for machine translation (Ranathunga et al., 2024) and general-purpose tasks (Reid and Artetxe, 2022a).
@@
Model Size: Recent studies indicate that scaling up language models enhances their capability to excel in diverse and complex reasoning tasks (Wei et al., 2022, 2023; Lu et al., 2023). Follow-up studies (Wei et al., 2023) further illustrate distinct behavioral differences between larger and smaller models.
@@ 3.1 Language
We use three criteria for language selection. Firstly, we select languages well covered by mLLMs as our goal is to assess how parallel data enhances performance after pre-training an mLLM with monolingual data. Secondly, the selected languages should be also covered by several different evaluation benchmarks, allowing for robust evaluation across diverse downstream tasks. Lastly, we select typologically diverse languages, enabling our investigation to generalize to a wide range of relatively low-resource languages. Thus, we select five languages: Arabic (ar), Spanish (es), Hindi (hi), Vietnamese (vi) and Chinese (zh).
@@ 3.3 Training
We select the BLOOM series (Scao et al., 2022) for our investigation due to its offering of different sizes of mLLMs which well cover the five target languages under consideration. [^2] Additional experiments with XGLM, as presented in §A, yield conclusions consistent with those of BLOOM. The pretraining data size for the five target languages ranges from 23GB (Hindi) to 452GB (English). We explore BLOOM models of various sizes, including 7B1, 3B, and 1B7. Due to limited computational resources, we use LoRA (Hu et al., 2022), which is known for its competitive performance compared to full-parameter training (Alves et al., 2023), for instruction tuning of BLOOM. We configure the learning rate to $1\times 10^{-4}$, weight decay to $0.1$, and set the rank of LoRA to $16$ based on preliminary experiment in §B. The maximum sequence length for both source and target sentences is set to $128$. To maintain consistency across experiments with different quantities of parallel corpora, we ensure a uniform training budget of 50K parallel sentences. Specifically, we calculate the number of epochs as 50K divided by the number of sentences considered from the OPUS100 dataset. The batch size is $128$, and we save the checkpoints every 20 steps.
Following Cahyawijaya et al. (2023), we construct the data for instruction tuning based on the parallel corpora by three distinct patterns: Machine Translation (MT), Translation Language Modeling (TLM), and Cross-Lingual Semantic Similarity (XSS). Table 1 presents the templates for these three objectives. Here, [SOURCE_LANG] and [TARGET_LANG] represent the language names of the source and target languages, respectively. In our study, we consider both English-to-target-language and target-language-to-English directions, where [SOURCE_LANG] represents English or [TARGET_LANG] represents English. For MT, [SOURCE_TEXT] and [TARGET_TEXT] refer to the parallel sentences in the source and target languages, respectively. For TLM, a portion of tokens in [TARGET_TEXT] are masked to generate [INPUT_TEXT]. For XSS, our objective is to predict whether parallel sentences [SOURCE_TEXT] and [TARGET_TEXT] are semantically similar, with [LABEL] being “Yes” or “No”. Specifically, we utilize the parallel corpora as positive examples and introduce perturbations to [TARGET_TEXT] to construct negative examples. We consider applying the objectives both individually and in combination. We tune the model with the objective of causal language modeling without loss mask.
@@ 3.4 Evaluation
To provide a thorough understanding of our evaluation procedures, we offer detailed prompts for each task in §C. In all experiments, we employ a 2-shot in-context learning approach, where the model is given two examples appended to the input query to aid in making predictions.
@@ 4.1 Quality of OPUS100
Relatively low-resource languages suffer more from low-quality issues. For relatively low-resource languages like Hindi, there are fewer high-quality parallel sentences compared to high-resource languages such as Spanish. Analysis of translation quality indicates that the English-Hindi pair has less than 20% of parallel sentences with high COMETWIKI scores (0.75-1), whereas the English-Spanish pair has around 45%. For sentence length, the English-Hindi pair contains 10% more short sentences (0-5 tokens) compared to high-resource language pairs. Moreover, both the English-Arabic and English-Hindi pairs exhibit about 10% more parallel sentences that may be in the wrong languages.
@@ 4.2 Effect of Quality
Parallel corpora containing noisy translations still improve results. Comparing the results of the experiment with $\tau_{c}=0$ (ID 1) to the original model (ID 0), there’s an average improvement of 0.4% for all tasks. The most notable improvements are observed in both bilingual tasks (en-xx) and in-language tasks. However, generating English for bilingual tasks yields degraded or marginally improved results. Experiment 1 exhibits 0.7% and 1.1% decrements in FLORES and MUSE respectively, with only a 0.3% improvement in MLQA.
@@ 5.1 Effect of Quantity Across Tasks
Table 4: Task performance (%) of BLOOM-7B1 after instruction tuning with the machine translation objective using varying amounts of parallel sentences, obtained with the best filtering strategy (ID 5) as shown in Table 3. |Sent| indicates the number of parallel sentences used for instruction tuning, with |Sent|=0 representing the original BLOOM-7B1 model.
@@
Using 10K parallel sentences leads to the optimal performance. The best overall performance is achieved with 10K parallel sentences, resulting in an average score of 54.0%. This setting yields the highest scores in MUSE and SIB. This aligns with the findings in Xu et al. (2023).
@@
Table 5: Language performance (%) of BLOOM-7B1 after instruction tuning with the machine translation objective using varying amounts of parallel sentences, obtained with the best filtering strategy (ID 5) as shown in Table 3. |Sent| indicates the number of parallel sentences used for instruction tuning, with |Sent|=0 representing the original BLOOM-7B1 model.
@@ 5.2 Effect of Quantity Across Languages
We delve deeper into the influence of parallel corpora quantity, as depicted in Table 5.
@@ 6 Objective
We explore the effectiveness of different objectives and their combinations, with results in Table 6.
@@ 8 Conclusion
Acknowledgements
This work was funded by DFG (SCHU 2246/14-1), the European Research Council (DECOLLAGE, ERC-2022-CoG #101088763), EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research, and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
Appendix A Additional Experiments on XGLM
We explore the effect of parallel corpora quality on XGLM-7.5B, and the results are shown in Table 8. As shown, experiments on XGLM-7.5B exhibits consistent findings as those on BLOOM-7B1.
ID Model Accuracy
0 XGLM-7.5B 52.8
$\tau_{c}$ $\tau_{l}$ LID
1 0 0 54.6
2 0.5 0 55.0
3 0.75 5 54.4
4 0.75 0 55.0
5 0.75 0 55.5
Table 8: Performance (%) of XGLM-7.5B on SIB after instruction tuning with the machine translation objective using 10K parallel corpora with various quality filtering strategies. Parameters include $\tau_{c}$ for COMETWIKI score threshold, $\tau_{l}$ for sentence length threshold, and LID indicating the adoption of language identification filtering.
Appendix B Effect of LoRA Rank
We conduct initial experiments on Arabic to explore the effect of setting the rank of LoRA, and the results are shown in Table 9. As shown, setting the rank of LoRA as 16 lead to best performance. Therefore, we use LoRA rank as 16 across all the experiments.
Rank Accuracy
8 59.0
16 62.0
32 60.0
128 61.0
256 60.5
512 59.5
Table 9: Effect of LoRA rank: Accuracy on SIB using English-Arabic parallel data to improve BLOOM-7B1.
Appendix C Prompt